OpenStudy (anonymous):

if u have a box 10cm in length and 7cm in width whats the perimeter and the area

OpenStudy (anonymous):

assuming that this box is only a 2d image(rectangle) not 3d then perimeter = 2(length) + 2(width) and area = length x width so perimeter = 2(10) + 2(7) = 20+14 = 34 area = 10(7) = 70

OpenStudy (anonymous):

To find the area of any object, you multiply the length by the width. In this case, 7cm x 10cm= 70cm^2. Just follow what @jayz657 has layed out. It's really simple.

OpenStudy (anonymous):

Also, don't forget that you must also multiply the units, in this case it's cm. So cm x cm =cm^2

OpenStudy (anonymous):

thank you so much just cant think todayhead is ful of mush